# ToDRE: Visual Token Pruning via Diversity and Task Awareness for Efficient Large Vision-Language Models

### v1

@@ Abstract
The representation of visual inputs of large vision-language models (LVLMs) usually involves substantially more tokens than that of textual inputs, leading to significant computational overhead. Several recent studies strive to mitigate this issue by either conducting token compression to prune redundant visual tokens or guiding them to bypass certain computational stages. While most existing work exploits token importance as the redundancy indicator, our study reveals that two largely neglected factors, namely, the diversity of retained visual tokens and their task relevance, often offer more robust criteria in token pruning. To this end, we design ToDRE, a two-stage and training-free token compression framework that achieves superior performance by pruning Tokens based on token Diversity and token-task RElevance. Instead of pruning redundant tokens, ToDRE introduces a greedy $k$-center algorithm to select and retain a small subset of diverse visual tokens after the vision encoder. Additionally, ToDRE addresses the “information migration” by further eliminating task-irrelevant visual tokens within the decoder of large language model (LLM). Extensive experiments show that ToDRE effectively reduces 90% of visual tokens after vision encoder and adaptively prunes all visual tokens within certain LLM’s decoder layers, leading to a 2.6$\times$ speed-up in total inference time while maintaining 95.1% of model performance and excellent compatibility with efficient attention operators.
@@ 1 Introduction
Building on the success of large language models (LLMs) [1, 3, 49, 51, 52], large vision-language models (LVLMs) [5, 22, 50, 56, 65] have recently emerged to handle a diverse range of multimodal understanding tasks by extending LLMs’ advanced reasoning capabilities. LVLMs typically convert visual inputs into visual tokens by a vision encoder (e.g., CLIP ViT-L/336 [45]), which are then aligned to the embedding space of LLM for multimodal interaction. Despite their impressive performance on various vision-language comprehension tasks, such as visual question answering [17, 19, 21, 41, 48] and video understanding [15, 27, 42, 55, 64], LVLMs incur prohibitive computational and memory costs during inference due to the massive number of visual tokens involved, significantly restricting their applicability in various practical tasks.
@@
Two representative approaches have recently been explored for improving the LVLM inference efficiency. The first is the model-centric approach. It addresses the problem by knowledge distillation [8], parameter quantization [57], and transformer replacement [44], but incurs significant computational resources in model retraining. The second is the data-centric approach. It works by token pruning [10, 35, 38, 46, 61] or block skipping [47], which has attracted increasing attention due to its training-free and architecture-agnostic nature. Beyond that, the data-centric approach offers an effective trade-off between inference efficiency and model performance and can serve as a complementary solution to the model-centric approach.
Figure 1: Compared with prior attention-based token pruning methods that overly rely on attention scores and may discard informative regions, our diversity-driven approach effectively preserves indispensable visual cues by retaining a small set of diverse visual tokens. This contributes to more robust token pruning and improves answer accuracy, as shown by the correct localization of the coffee cup in the example.
@@
Figure 2: We plot the ratio of text-to-visual (orange) and visual-to-text (blue) attentions in each LLM decoder layer. A clear “information migration” phenomenon can be observed: visual-to-text attention dominates in early layers, indicating massive absorption of visual information by text tokens, while both cross-modal attentions vanish in deeper layers, suggesting a shift toward unimodal text reasoning.
Figure 3: Performance of various token compression methods on LLaVA-NeXT-7B [37] across eight image-language comprehension benchmarks under a 90% reduction ratio. ToDRE outperforms other methods across all evaluated tasks.
@@
Existing token compression techniques typically remove visual tokens according to a predefined “redundancy” metric. This redundancy metric is often estimated by the attention scores between the visual tokens and other token(s) (e.g., text tokens [10, 62], or [CLS] token [46, 61]), an observed divergence of LLM’s outputs before and after token pruning [35, 59], or token similarity [6, 23, 63]. However, each above-mentioned criterion has clear shortcomings. Specifically, attention scores exhibit positional bias, such that tokens spatially close to the last token tend to receive higher weights [54]. Consequently, some informative tokens are erroneously discarded (Figure 2). Divergence-based methods require a held-out calibration set and model-specific distribution matching, which hinders quick adaptation to new LVLM backbones [35]. Similarity-based approaches instead involve merging redundant visual tokens into retained ones with highest similarity, which has been validated to be less effective than direct token pruning [20]. Beyond the above issues, we observe an “information migration” phenomenon (Figure 2), where visual information is progressively transferred to subsequent text tokens within the first half of the LLM-decoder layers. However, most prevailing work [10, 46, 61, 62] passes the entire set of visual tokens from the prefilling stage into decoding, thereby incurring unnecessary computations.
To more accurately quantify the redundancy of visual tokens and accelerate LVLM inference, we propose ToDRE, a simple yet effective token pruning method that jointly considers token diversity among visual tokens and token relevance with target tasks. Prior to the LLM’s decoder, ToDRE employs a greedy k-center algorithm to select a subset of visual tokens, where at each step the token with the greatest minimum distance to the current subset is added. This token selection strategy circumvents the positional bias introduced by textual attention, thereby preserving a broader spectrum of visual information and making the token representation more robust at high pruning ratios. To combat the “information migration” effect observed in the LLM’s decoder, ToDRE adaptively selects a layer in the latter half—where cross-modal attention has significantly diminished—and drops all visual tokens within that layer. As a result, our relevance-guided pruning approach enables continuous inference-time efficiency gains as output length increases.
In summary, our major contributions are threefold:
  • Rethink the Criterion of Token Redundancy. We provide a systematic analysis showing that existing attention-based redundancy indicators are sub-optimal. By collectively examining inter-token diversity and token-task relevance, we uncover two crucial yet orthogonal factors previously overlooked, enabling more effective token pruning.
@@
  • Identify Token Diversity and Token-task Relevance as Key Factors. Motivated by the insight “stop eliminating redundant tokens—start retaining diverse ones”, we introduce a two-stage plug-and-play token compression technique that is fully compatible with efficient attention operators [13] without requiring any additional training.
  • Superior Performance with Extreme Compression. We conduct comprehensive experiments over two model sizes and twelve image- and video-language benchmarks, demonstrating the notable superiority of the proposed ToDRE. While existing one-stage methods [10, 39, 61] prune up to 90% of visual tokens, our approach further eliminates all remaining visual tokens in an additional stage to obtain higher efficiency, yet still outperforms the second-best method by 2.4% (95.1% vs. 92.7%) on LLaVA-NeXT-7B [37].
@@ 2.1 Large Vision-Language Models
Large vision-language models (LVLMs) [5, 50, 65] have demonstrated remarkable advancements by extending the reasoning capabilities of pretrained LLMs [3, 51, 52] to image and video comprehension tasks. Typically, LVLMs employ a vision encoder to extract visual features, which are subsequently projected into the LLM’s embedding space via a visual projector (e.g., Q-Former [26] or MLP [30, 37]). To process real-world high-resolution images, previous LVLMs [4, 36] resize input images to a fixed resolution, which introduces geometric distortion and degrade fine-grained local details. To tackle this, subsequent studies adopt dynamic tiling [11, 25, 37], which partitions images into regions and encodes each region independently using a shared vision encoder. However, dynamic tiling can yield thousands of visual tokens, significantly increasing computational overhead. This issue becomes even more pressing in video-based LVLMs [5, 33], since processing multiple video frames demands significantly more visual tokens. These challenges highlight the urgent need for accelerating LVLM inference in resource-constrained real-world environments.
@@ 2.2 Token Compression for LVLMs
Given that spatially redundant visual tokens outnumber information-dense text tokens by tens to hundreds of times [43], one natural solution to optimize LVLM inference is visual token compression. Several earliest attempts [7, 28, 29, 58] modify model components and introduce additional training costs. More recently, training-free token compression methods have been widely adopted due to their efficiency and effectiveness. These methods can be categorized into two main groups: (1) Token compression in the vision encoder [6, 32, 46], the LLM decoder [10, 35, 62], or both [20]: For example, ToMe [6] reduces tokens in the vision encoding phase by merging redundant tokens via a binary soft-matching algorithm. Other approaches prune tokens during the LLM decoding stage by evaluating token redundancy through criteria such as attention scores with text tokens [10, 62] or observed divergence with LLM outputs [35, 59]. Subsequent studies [20, 38, 66] perform token compression during both stages to further enhance inference efficiency. (2) Token compression in LLM embedding space [2, 61]: FasterVLM [61] can be a representative, which measures the token redundancy more accurately by the cross-attentions between the [CLS] token and visual tokens. Unlike previous methods, our proposed ToDRE simultaneously reduces tokens in both the LLM’s embedding space and the LLM decoder. Our two-stage approach effectively captures both token diversity and task relevance—two critical aspects previously overlooked—achieving superior inference efficiency while maintaining competitive performance.
@@ 3 Preliminary Analysis
Recently, abundant visual token compression techniques have emerged. Most approaches [2, 10, 35, 54, 61] reduce computational redundancy only within partial stages of LVLM inference process, lacking a systematic analysis and overall consideration. To bridge this gap, we provide a deeper analysis organized as follows. In Section 3.1, we review the fundamental architecture and processing flow of existing LVLMs, identifying where redundant computation arises. In the following Section 3.2, we further provide empirical observations and examine the limitations of existing redundancy-reduction strategies, which motivate us to propose a two-stage token pruning method. In Appendix B a theoretical proof is presented to validate the underlying rationale and structural integrity of the proposed two-stage paradigm.
3.1 Identifying Redundancy: Redundant Computations in LVLM Processing Pipeline
@@
Typically, existing LVLMs consist of three main components: a vision encoder, a vision-language projector, and a LLM decoder. Both the encoder and decoder are built upon the Transformer blocks [53]. Given a visual input $V$, the vision encoder extracts visual features, which are then mapped into a sequence of visual token embeddings $E_{v}$ by the vision-language projector, aligned with the LLM’s textual embedding space. Then, $E_{v}$ is concatenated with text embeddings $E_{t}$ and system prompt embeddings $E_{s}$ to form the input sequence for LLM. During the LLM’s prefilling stage, all input tokens interact via self-attention to generate a contextualized representation, denoted as $X=\{\bm{z}_{s_{1}},\dots,\bm{z}_{s_{L}},\bm{z}_{v_{1}},\dots,\bm{z}_{v_{M}},\bm{z}_{t_{1}},\dots,\bm{z}_{t_{N}}\}$, where $L$, $M$ and $N$ denote the sequence lengths of system prompt token $\bm{Z}_{s}$, visual token $\bm{Z}_{v}$, and text token $\bm{Z}_{t}$, respectively. At each Transformer layer, $X$ is projected into keys and values, which are then stored as KV cache. In the subsequent decoding stage, keys and values are computed and added only for newly generated tokens, while previously computed key-value pairs are retrieved from the cache directly.
@@
Prior studies [20, 38] have shown that the dominant contributors to inference cost in LVLMs are the vision-encoding stage, the LLM’s prefilling stage, and the LLM’s decoding stage, each of which incurs substantial self-attention and feed-forward network (FFN) computations. Following previous studies [10, 54], we formulate the calculation of floating-point operations (FLOPs) as follows:
@@
where $T$ is the number of transformer layers; $n$ and $L$ respectively denote the lengths of the input and output sequences; $d$ is size of the hidden state; and $m$ is the intermediate dimension of the FFN. We take LLaVA-NeXT-7B [37], which employs CLIP-ViT-Large-Patch14 [45] vision encoder and Vicuna-7B-v1.5 [12] LLM decoder, as an example. The relative ratio of FLOPs (with $n$=3000 and $L$=20) is approximately encoding:prefilling:decoding $\approx$ 1:63.6:0.4. When scaled to LLaVA-NeXT-13B, the relative ratio shifts to 1:121.1:0.8, indicating that the LLM’s prefilling and decoding stages roughly double their share of the total computational cost. This underscores the importance of pruning visual tokens as early as possible—ideally prior to or during the LLM prefilling stage—to mitigate the exploding computational burden.
3.2 Rethinking Redundancy: Both Intra- and Cross-Modal Redundancies Matter
@@
A common practice is to identify the most “important” tokens based on predefined criteria, and then apply token-level pruning or merging strategies. Attention-based methods—such as averaging attention scores [10] or leveraging attention from the [CLS] token to visual tokens [61]—are widely adopted. However, such methods suffer from attention shift, where causal decoding biases attention toward later-positioned visual tokens [54]. Moreover, attention distributions are often imbalanced: [CLS]-based attention is overly concentrated, while text-to-visual attention tends to be dispersed and noisy [61]. These limitations motivate a natural rethinking: what is the essence of visual token redundancy? While earlier studies have not delved deeply into this issue, we argue that token redundancy manifests in two orthogonal components: intra-modal redundancy within the visual signal, and cross-modal redundancy between visual and textual modalities.
@@
Intra-modal redundancy occurs when visual tokens exhibit significant similarity, since highly similar tokens contribute little unique information and are thus redundant. Such redundancy can be identified using visual-only signals, typically by measuring cosine similarity. Then, the problem reduces to selecting a minimally redundant subset of tokens. Here, instead of relying on complex designs for redundancy detection, we find that retaining a maximally diverse set of tokens more effectively preserves the visual representation (see Figure 2). This observation motivates us to introduce the Diversity-driven Visual Token Selection, acting as the first stage of ToDRE prior to LLM prefilling.
@@
On the other hand, LVLM’s multimodal comprehension heavily depends on textual cues [60], giving rise to cross-modal redundancy where visual tokens that are less relevant to the textual information can be safely pruned. In this view, the attention scores between visual and text modalities during the LLM prefilling stage offer a simple yet reliable signal for token reduction. By treating cross-modal attention as a unified whole, we avoid the previously mentioned limitations of attention-based selection strategies. Building on the concept of decoding-stage information migration proposed in VTW [35], we further analyze its behavior during the LLM prefilling stage. As shown in Figure 2, cross-modal attention is prominent in early layers and gradually diminishes in deeper layers, revealing the information migration phenomenon during prefilling: early layers prioritize cross-modal interaction, while deeper layers focus primarily on uni-modality processing. This finding drives us to propose the Relevance-driven Visual Token Reduction, serving as the second stage of ToDRE during LLM prefilling.
4 ToDRE: Two-stage Visual Token Pruning via Diversity and Task Awareness

### v2

# ToDRE: Effective Visual Token Pruning via Token Diversity and Task Relevance

@@ Abstract
[…]
Figure 1: (a–c): Different from the prevalent visual token pruning approach [chen2024fastv, zhang2024fastervlm] that overly relies on attention scores, the proposed ToDRE incorporates token diversity and task relevance, two largely neglected yet critical factors that help preserve indispensable and informative visual cues and improve pruning robustness and answer accuracy as illustrated in the coffee cup localization task. (d): Quantitative experiments over eight image-language comprehension benchmarks demonstrate the superior and consistent effectiveness of our proposed ToDRE.
@@ 1 Introduction
Leveraging the superior reasoning capability of large language models (LLMs) [achiam2023gpt, bai2023qwen, team2023gemini, touvron2023llama1, touvron2023llama2], large vision-language models (LVLMs) [bai2025qwen25vl, hurst2024gpt4o, team2025kimivl, wu2024deepseekvl2, zhu2025internvl3] have achieved impressive performance in various multimodal understanding tasks such as visual question answering [goyal2017vqav2, gurari2018vizwiz, hudson2019gqa, lu2022sqa, singh2019textvqa] and video understanding [fu2024videomme, li2024mvbench, mangalam2023egoschema, wu2024longvideobench, zhou2024mlvu]. LVLMs convert visual inputs into visual tokens and align the converted visual tokens with text tokens for various multimodal understanding tasks. However, the inference of LVLMs often incurs prohibitive computational and memory costs due to the massive number of visual tokens involved, significantly restricting LVLM applicability in various downstream tasks.
@@
Two representative approaches have recently been explored for improving the LVLM inference efficiency. The first approach is model-centric. It speeds up the inference via knowledge distillation [cai2024llavakd], parameter quantization [xie2024quant], or transformer replacement [qiao2024vlmamba]. However, this approach requires model retraining which incurs significant computational resources. The second approach is data-centric. It works by token pruning [chen2024fastv, lin2025vtw, liu2024mustdrop, shang2024prumerge, zhang2024fastervlm] or block skipping [shukor2024skip], and has attracted increasing attention due to its training-free and architecture-agnostic nature. Besides, the data-centric approach strikes a great balance between the inference efficiency and the model performance, offering a complementary solution to the model-centric approach.
@@
Figure 2: Text-to-visual attention (blue) and visual-to-text attention (orange) in each LLM decoder layer. We observe a clear pattern of “information migration”: cross-modal attention (both visual-to-text and text-to-visual) is high in early layers, reflecting active information exchange, but gradually diminishes in deeper layers as the model shifts toward unimodal text reasoning.
@@
Most existing token pruning techniques compress visual tokens by estimating “redundancy” from a single metric, such as cross-modal attention between visual and other-modality tokens [chen2024fastv, zhang2024sparsevlm, shang2024prumerge, zhang2024fastervlm], visual token similarity [bolya2022tome, jiang2025gprune, zhao2024gsearch], or the divergence of LLM’s outputs before and after token pruning [lin2025vtw, ye2025fitprune]. However, attention scores exhibit clear positional bias [wen2025dart] that tends to discard informative tokens erroneously (Figure˜1 (b)). Similarity-based approach merges similar visual tokens whose performance is often clearly lower than direct token pruning [han2024ficoco]. Using output divergence requires a held-out calibration set and model-specific distribution matching, hindering quick adaptation towards new LVLM backbones [lin2025vtw]. Beyond the above issues, we observe an “information migration” phenomenon (Figure˜2): cross-modal attention (both visual-to-text and text-to-visual) is strong in early layers but fades in deeper layers, suggesting that visual information is progressively absorbed into text representations within the first half of the LLM decoder. Given that output tokens exhibit near-zero attention to visual tokens during decoding (see Appendix), most existing work [chen2024fastv, shang2024prumerge, zhang2024fastervlm, zhang2024sparsevlm] passes all remaining visual tokens from the prefilling stage into decoding, thereby incurring unnecessary computations.
We design ToDRE, a simple yet effective token pruning technique that incorporates both visual token diversity and task-specific token relevance for effective token pruning and efficient LVLM inference. ToDRE performs token pruning in the embedding space prior to LLM input and during the LLM prefilling stage. First, we introduce a greedy max-sum diversification algorithm that iteratively identifies and preserves visual tokens that have minimal cumulative similarity to the selected tokens. Such token selection in LLM embedding space circumvents the positional bias introduced by attention-based metrics, thereby preserving a broad spectrum of visual information and enhancing the token representativeness at high pruning ratios. In addition, ToDRE leverages the “information migration” mechanism by adaptively selecting one layer in the latter half of the LLM decoder (where cross-modal attention has significantly diminished) and drops all visual tokens within that layer. This layer-level pruning removes visual tokens irrelevant to the given question and thus further eliminates redundant computation during inference. As a result, this relevance–guided pruning enables continuous inference-time efficiency gains as the decoding length increases. As shown in Figure˜1 (c–d), ToDRE’s two-stage design enables effective visual token compression while preserving unique visual information and maintaining strong accuracy.
In summary, our major contributions of this work are threefold:
  • Revisit redundancy indicators. First, we re-examine the principles of existing indicators on token redundancy and identify their constraints via systematic and comprehensive analysis. On top of that, we prove that inter-token diversity and token-task relevance are two orthogonal factors, and treating them separately enables more effective token pruning.
@@
  • Propose a training-free and plug-and-play framework. Second, we design a two-stage plug-and-play token pruning technique that is fully compatible with efficient attention operators [dao2022flashattention] without requiring any additional training.
  • Conduct extensive empirical validation. Third, extensive experiments over four widely adopted LVLMs and twelve multimodal benchmarks show that ToDRE achieves superior token pruning consistently.
@@ 2.1 Large Vision-Language Models
Large vision-language models (LVLMs) [bai2025qwen25vl, team2025kimivl, zhu2025internvl3] have demonstrated remarkable advancements by extending the reasoning capabilities of pretrained LLMs [bai2023qwen, touvron2023llama1, touvron2023llama2] to image and video comprehension tasks. Typically, LVLMs employ a vision encoder to extract visual features, which are subsequently projected into the LLM’s embedding space via a visual projector (e.g., Q-Former [li2023blip] or MLP [li2024mini, liu2024llavanext]). To process real-world high-resolution images, previous LVLMs [bai2023qwenvl, liu2024llava15] resize input images to a fixed resolution, which introduces geometric distortion and degrades fine-grained local details. To tackle this, subsequent studies adopt dynamic tiling [chen2024far, li2024llavaonevision, liu2024llavanext], which partitions images into regions and encodes each region independently using a shared vision encoder. However, dynamic tiling can yield thousands of visual tokens, significantly increasing computational overhead. This issue becomes even more pressing in video-based LVLMs [bai2025qwen25vl, lin2023videollava], since processing multiple video frames demands significantly more visual tokens. These challenges highlight the urgent need for accelerating LVLM inference in resource-constrained real-world environments.
@@ 2.2 Token Compression for LVLMs
Given that spatially redundant visual tokens outnumber information-dense text tokens by tens to hundreds of times [marr2010vision], one natural solution to optimize LVLM inference is visual token compression. Several earliest attempts [cai2024matryoshka, li2024tokenpacker, li2024llamavid, yao2024deco] modify model components and introduce additional training costs. More recently, training-free token compression methods have been widely adopted due to their efficiency and effectiveness. These methods can be categorized into two main groups: (1) Token compression in the vision encoder [bolya2022tome, liang2022evit, shang2024prumerge], the LLM decoder [chen2024fastv, lin2025vtw, zhang2024sparsevlm], or both [han2024ficoco]: For example, ToMe [bolya2022tome] reduces tokens in the vision encoding phase by merging redundant tokens via a binary soft-matching algorithm. Other approaches prune tokens during the LLM decoding stage by evaluating token redundancy through criteria such as attention scores with text tokens [chen2024fastv, zhang2024sparsevlm] or observed divergence with LLM outputs [lin2025vtw, ye2025fitprune]. Subsequent studies [han2024ficoco, liu2024mustdrop, zhu2024focusllava] perform token compression during both stages to further enhance inference efficiency. (2) Token compression in LLM embedding space [alvar2025divprune, zhang2024fastervlm]: A representative example is FasterVLM [zhang2024fastervlm], which measures the token redundancy more accurately by the cross-attentions between the [CLS] token and visual tokens. Unlike previous methods, our proposed ToDRE simultaneously reduces tokens in both the LLM embedding space and the LLM decoder. Our two-stage approach effectively captures both visual token diversity and token-task relevance—two orthogonal yet critical aspects previously overlooked—achieving superior inference efficiency while maintaining competitive performance.
@@ 3 Preliminary Analysis
Recently, numerous visual token compression techniques have emerged. Most approaches [alvar2025divprune, chen2024fastv, lin2025vtw, wen2025dart, zhang2024fastervlm] reduce computational redundancy only within partial stages of the LVLM inference process, lacking a systematic analysis and overall consideration. To bridge this gap, we provide a deeper analysis organized as follows. In Section˜3.1, we review the fundamental architecture and processing flow of existing LVLMs, identifying where redundant computation arises. In the following Section˜3.2, we further provide empirical observations and examine the limitations of existing redundancy-reduction strategies, which motivate us to propose a two-stage token pruning method. In the Appendix, a theoretical proof is presented to validate the underlying rationale and structural integrity of the proposed two-stage paradigm.
3.1 Computational Overhead in LVLM Processing Pipeline
@@
Typically, existing LVLMs consist of three main components: a vision encoder, a vision-language projector, and a LLM decoder. Both the encoder and decoder are built upon the Transformer blocks [vaswani2017transformer]. Given a visual input $V$, the vision encoder extracts visual features, which are then mapped into a sequence of visual token embeddings $E_{v}$ by the vision-language projector, aligned with the LLM textual embedding space. Then, $E_{v}$ is concatenated with text embeddings $E_{t}$ and system prompt embeddings $E_{s}$ to form the input sequence for LLM. During the LLM’s prefilling stage, all input tokens interact via self-attention to generate a contextualized representation, denoted as $X=\{\bm{z}_{s_{1}},\dots,\bm{z}_{s_{L}},\bm{z}_{v_{1}},\dots,\bm{z}_{v_{M}},\bm{z}_{t_{1}},\dots,\bm{z}_{t_{N}}\}$, where $L$, $M$ and $N$ denote the sequence lengths of system prompt token $\bm{Z}_{s}$, visual token $\bm{Z}_{v}$, and text token $\bm{Z}_{t}$, respectively. At each Transformer layer, $X$ is projected into keys and values, which are then stored as KV cache. In the subsequent decoding stage, keys and values are computed and added only for newly generated tokens, while previously computed key-value pairs are retrieved from the cache directly.
@@
Prior studies [han2024ficoco, liu2024mustdrop] have shown that the dominant contributors to inference cost in LVLMs are the vision-encoding stage, the LLM prefilling stage, and the LLM decoding stage, each of which incurs substantial self-attention and feed-forward network (FFN) computations. Following previous studies [chen2024fastv, wen2025dart], we formulate the calculation of floating-point operations (FLOPs) as follows:
@@
where $T$ is the number of transformer layers; $n$ and $L$ respectively denote the lengths of the input and output sequences; $d$ is size of the hidden state; and $m$ is the intermediate dimension of the FFN. We take LLaVA-NeXT-7B [liu2024llavanext], which employs CLIP-ViT-Large-Patch14 [radford2021clip] vision encoder and Vicuna-7B-v1.5 [chiang2023vicuna] LLM decoder, as an example. The relative ratio of FLOPs (with nn=3000 and LL=20) is approximately encoding:prefilling:decoding $\approx$ 1:63.6:0.4. When scaled to LLaVA-NeXT-13B, the relative ratio shifts to 1:121.1:0.8, indicating that the LLM’s prefilling and decoding stages roughly double their share of the total computational cost. This underscores the importance of pruning visual tokens as early as possible—ideally prior to or during the LLM prefilling stage—to mitigate the exploding computational burden.
3.2 Intra- and Inter-Modal Redundancy
@@
A common practice is to identify the most “important” tokens based on predefined criteria, and then apply token-level pruning or merging strategies. Attention-based methods—such as averaging attention scores [chen2024fastv] or leveraging attention from the [CLS] token to visual tokens [zhang2024fastervlm]—are widely adopted. However, such methods suffer from attention shift, where causal decoding biases attention toward later-positioned visual tokens [wen2025dart]. Moreover, attention distributions are often imbalanced: [CLS]-based attention is overly concentrated, while text-to-visual attention tends to be dispersed and noisy [zhang2024fastervlm]. These limitations motivate a natural rethinking: what is the essence of visual token redundancy? While earlier studies have not delved deeply into this issue, we argue that token redundancy manifests in two orthogonal components: intra-modal redundancy within the visual signal, and cross-modal redundancy between visual and textual modalities.
@@
Intra-modal redundancy occurs when visual tokens exhibit significant similarity, since highly similar tokens contribute little unique information and are thus redundant. Such redundancy can be identified using visual-only signals, typically by measuring cosine similarity. Then, the problem reduces to selecting a minimally redundant subset of tokens. Here, instead of relying on complex designs for redundancy detection, we find that retaining a maximally diverse set of tokens more effectively preserves the visual representation. This observation motivates us to introduce the Diversity-driven Visual Token Selection, acting as the first stage of ToDRE prior to LLM prefilling.
@@
On the other hand, LVLM’s multimodal comprehension heavily depends on textual cues [zhang2023visionlanguagemodelsgoodguesser], giving rise to cross-modal redundancy where visual tokens that are less relevant to the textual information can be safely pruned. In this view, the attention scores between visual and text modalities during the LLM prefilling stage offer a simple yet reliable signal for token reduction. By treating cross-modal attention as a unified whole, we avoid the previously mentioned limitations of attention-based selection strategies. Building on the concept of decoding-stage information migration proposed in VTW [lin2025vtw], we further analyze its behavior during the LLM prefilling stage. As shown in Figure˜2, cross-modal attention is prominent in early layers and gradually diminishes in deeper layers, revealing the information migration phenomenon during prefilling: early layers prioritize cross-modal interaction, while deeper layers focus primarily on uni-modality processing. This finding drives us to propose the Relevance-driven Visual Token Reduction, serving as the second stage of ToDRE during LLM prefilling.
4 Visual Token Pruning with Token Diversity and Task Relevance